OpenStudy (anonymous):

Simplify: (5x2 + 3x + 4) + (2x2 − 6x + 3) 7xsq. + 9x + 7 7xsq. + 3x + 7 7xsq.− 3x + 1 7xsq. − 3x + 7

OpenStudy (noseboy908):

think of it like this: 3x-6x=?, and 4+3=?.

OpenStudy (anonymous):

answer?

OpenStudy (anonymous):

the 2 after the x's is sq

OpenStudy (noseboy908):

yeah, I got that. First answer my previous two questions: 3x-6x=?, and 4+3=?.

OpenStudy (anonymous):

-3 7

OpenStudy (noseboy908):

-3 what? don't forget the x, but otherwise correct. so the end answer would be...

OpenStudy (anonymous):

Is the answer D.

OpenStudy (noseboy908):

correct!